find k so that the line through 4,k and 2,-1 is perpendicular to y= 2.5x-3
Answer by mananth(16946)

line through 4,k and 2,-1 is perpendicular to
y= 2.5x-3
slope = 2.5=>5/2
perpendicular line will have slope -2/5 ( negative reciprocal)
slope of line between two points
(y2-y1)/(x2-x1)=m
(-1-k)/(2-4)= -2/5
-1-k=-2*-2/5
-1-k=-4/5
k=-1+4/5
k=-1/5
k=-0.2
m.ananth@hotmail.ca